OpenStudy (anonymous):

Eight tickets to a play cost $96. How much do 3 tickets cost? A. $12 B. $36 C. $48 D. $60

OpenStudy (anonymous):

36

OpenStudy (31356):

Set up a ratio

OpenStudy (31356):

B]36

OpenStudy (31356):

Multiply 96*3 Then divide the answer by 8 and then you get 36

OpenStudy (31356):

Medal?

OpenStudy (anonymous):

So you set up a ratio. \[\frac{ $96 }{ 8 tickets }\]

OpenStudy (anonymous):

Then you divide 96 by 8 to get the cost of 1 ticket. 96/8=12 so \[\frac{ $12 }{ 1 ticket }\]

OpenStudy (31356):

Or you can set it up 8/96 It doesn't matter at least the labels are the same Dollars to tickets or Ticket to dollars

OpenStudy (31356):

Hope that helps!

OpenStudy (anonymous):

You now multiply 12 by 3 since you are looking for the price of 3 tickets. 12*3=36 so now you finally get B. $36 or \[\frac{ $36 }{ 3 tickets }\]